OpenStudy (julierosehead):

Simplify the following expressions Please help these are the last two I have to solve. a. 5-2(x-3) b. 5(x-z) + 4(y-x)

OpenStudy (princeharryyy):

a) 5-2x+6 = -2x+11

OpenStudy (princeharryyy):

b) 5x-5z+4y-4x = x-5z+4y

OpenStudy (julierosehead):

in equation a, where did 6 come from?!

OpenStudy (princeharryyy):